# pulsars; FRBs

### 08 Apr 2024 v1

@@ 5 Discussion and conclusions
We have improved the solutions of the pulsar equation when the tip of the closed-line region, the so-called Y-point, approaches the light cylinder. The new solutions were obtained under the assumption that the separatrix surface between open and closed field lines is a zero-thickness contact discontinuity. The new solutions are similar to the old ones of Contopoulos et al. (1999); Contopoulos (2005); Timokhin (2006) when the Y-point lies inside about $80\%$ of the light cylinder radius. As the Y-point approaches closer to the light cylinder however, the solutions diverge from the standard ones found in the literature, and in the extreme limit that the closed-line region touches the light cylinder, the open-line region disappears completely. As is seen in eq. (12) in Appendix 2 and in figure 4, the closer the Y-point to the light cylinder, the stronger the divergence of $B_{p}$ right inside the Y-point at the tip of the closed-line region needs to be. We see no other way to implement this divergence of $B_{p}(x_{\rm Y}\rightarrow 1^{-})$ except with more and more field lines to enter the closed-line region and compress the magnetic field at its tip. This effect is clearly seen in the sequence of solutions shown in figure 3. In the limit where $B_{p}\rightarrow\infty$ as $x_{\rm Y}=1^{-}$, all field lines enter the closed-line region confined inside the light cylinder, and the open line region disappears completely (figure 1).
@@
One may legitimately ask why these solutions were never before obtained in the literature. There are several reasons. Firstly, in all previous solutions, the separatrix surface along which the magnetospheric electric curent circuit closes on the star had non-zero thickness due to the finite resolution of the code. In particular, at its tip, the outer part of the separatrix forms a Y-point that touches the light cylinder, while the inner part forms a clear T-point some distance inside it. This is clearly explained in figure 4 of Contopoulos et al. (2024). Unfortunately, the ideal force-free conditions of eq. (2) are not valid inside current sheets (as e.g. in figure 1b of Spitkovsky (2006)). Nevertheless, the fact that PIC simulations also form a separatrix that contains a nonzero amount of poloidal magnetic flux (see high resolution detail in figure 9 of Hu and Beloborodov (2022)) led us to consider the possibility that this effect is indeed physical. It is interesting that the separatrix current sheet contains a nonzero guide field and is thus different from the equatorial current sheet that is an ordinary Harris-type current sheet with $180^{\circ}$ field inversion across it. It also seems that the width of the former type of current sheet grows with time as is seen in figure 1c of Spitkovsky (2006). This effect merits further investigation with hybrid PIC-FFE simulations (Soudais, Cerutti & Contopoulos 2024, in preparation). Now that the separatrix has zero thickness, it forms a clear T-point at its tip which indeed lies some distance inside the light cylinder. Secondly, time dependent simulations always start from a vacuum dipole configuration that extends to all space and therefore, some fraction of it already crosses the light cylinder. After the central star is set into rotation, the magnetosphere is divided into open and closed field lines, i.e. there is no chance to obtain a configuration totally enclosed inside the light cylinder as the solution shown in figure 1. And thirdly, after the time-dependent evolution relaxes to a steady-state solution, it is reasonable to expect that it will further relax to the lowest energy solution. Therefore, time-dependent simulations cannot yield the full sequence of solutions up to the latter singular solution as our Machine Learning methodology is able to do. It is interesting that, now that the full sequence of solutions is obtained, it is shown that the solution where $x_{\rm Y}\sim 80\%\ R_{\rm LC}$ is a minimum energy solution, and this is why previous time dependent high-resolution PIC simulations relax to that solution. We have thus improved upon the result of Contopoulos et al. (2024) which placed the minimum energy solution around $x_{\rm Y}=92\%\ R_{\rm LC}$.
@@
We argue that the confined singular solution of figure 1 and all solutions between that one and the minimum energy solution are unstable, thus if the magnetospheric configuration of figure 1 forms during a supernova explosion or during a high accretion event from a surrounding disk or wind from a donor star, it will transition to the minimum energy solution by emitting a burst of energy equal to the energy difference between these two solutions. We propose that this burst may be related to an FRB. We would like to notice here that, as long as the surrounding material is there, the confined solution cannot transition to a standard solution in which part of the magnetosphere extends to infinity, and thus the accumulated electromagnetic energy of the confined solution cannot be released. What happens when the surrounding material dissolves, and how fast electromagnetic energy is released remains to be seen in a time-dependent numerical simulation of this transition. In case the pulsar magnetosphere undergoes a later accretion event from an external disk or wind from a donor star that will re-confine it to the solution of figure 1, the FRB will repeat. We must acknowledge of course that a few FRBs have already been observed from a magnetar, and that the cosmological sky rate of FRBs exceeds that of core-collapse supernovae, therefore, our model may just be a rare kind of FRB, if feasible at all. Moreover, we haven’t answered why would the emission of this state transition be in the MHz to GHz radio regime, and how it would escape the very dense accretion environment that formed the confined solution of figure 1. Our model certainly merits further investigation with detailed MHD numerical simulations.
@@
What is also interesting to investigate is whether a confined solution exists for non-aligned (oblique) pulsars in which $\partial{\bf B}/\partial t\neq{\bf 0}$. We expect that oblique rotators would radiate electromagnetic radiation to infinity, thus, magnetic field lines would extend to infinity. Therefore, it seems impossible for an oblique magnetosphere to be totally confined inside the light cylinder. It may thus be that the confined solution of the type shown in figure 11 is only possible in aligned pulsars.
@@
Conceptualization and supervision, I.C.; methodology, I.C., K.N.G.; software and visualization, D.N., I.D.; writing—original draft preparation, I.C.; writing—review and editing, all authors. All authors have read and agreed to the published version of the manuscript.
@@
This research received no external funding.
@@ Acknowledgements.
Appendix 1
In the first paper of this series (Dimitropoulos et al., 2024, Paper I), a new methodology was presented for solving the pulsar equation which essentially consists of the following steps:
@@
    It is very helpfull to deal with the equatorial current sheet in the open line region by mathematically reversing the polarity of the magnetic field that emanates from the south pole of the star. By doing this, the mathematical discontinuity in the equatorial region disappears, thus one do not need to worry about it computationally. At the end of the calculation, we will return to the original polarity of the magnetic field, and the equatorial current sheet will reappear.
@@
    The shape $r_{\rm S}(\theta)$ of the separatrix is readjusted so that continuity of $B^{2}-E^{2}$ is achieved at all points across it. One way to do that is to move the separatrix radially at each point with an amount that is proportional to the difference $(B^{2}-E^{2})_{\rm IN}-(B^{2}-E^{2})_{\rm OUT}$.
@@
    Repeat the procedure till pressure balance is achieved over the whole separatrix and its shape stabilizes.
@@
As it is acknowledged in Dimitropoulos et al. (2024), the numerical accuracy of our method is not on par yet with that of standard methods such as finite difference, finite volume or spectral methods. This is also the case in the solution of Stefanou et al. (2023). On the other hand, the numerical treatment of current sheets by standard methods it too is problematic (numerical current sheets have unphysical thicknesses and the numerical dissipation there is questionable). In order to improve our results in future iterations of this work, we will try to combine standard methods outside current sheets with our methodology of moving the separatrix. This is still work in progress. Modulo the above disclaimers, we applied this method to obtain the sequence of solutions shown in figure 3.
Appendix 2
We discuss here the shape of the tip of the closed-line region, the so-called Y-point. In Paper I it was found that
@@
    The Y-point is a clear T-point as predicted by Uzdensky (2003), not a Y-point as is seen in all numerical solutions of the axisymmetric pulsar magnetosphere to-date, except for Contopoulos et al. (2024). This is probably due to the fact that in all previous numerical simulations, the separatrix between closed and open field lines had an unphysically large thickness (sometimes on the order of the light cylinder!), thus it was not a true contact discontinuity as in Paper I.
@@
Both of the above results modify our understanding of how fast pulsars spin down. In the standard solution, the two regions are separated by a separatrix surface that contains a charged current sheet. One very interesting point in understanding the operation of the pulsar magnetosphere is how close to the light cylinder can the Y-point lie. According to eq. (3), the pressure balance condition across the separatrix current sheet may be written as
$\left.B_{p}^{2}(1-x^{2})\right|_{\rm IN}=\left.B_{p}^{2}(1-x^{2})\right|_{\rm OUT}+\frac{I(\Psi_{\rm S})^{2}}{x^{2}}$ (12)
Here, $I(\Psi_{\rm S})$ is the electric return current flowing along the separatrix, and $p$-indices denote poloidal ($r,\theta$) components of the fields. Let us first repeat the main argument why the tip of the closed line region on the equator is a T-point. The poloidal magnetic field $\left.B_{p}\right|_{\rm OUT}$ in the right hand side of eq. (12) is obviously equal to zero right above and below the equatorial current sheet. The third term, however, is finite because of the presence of the nonzero equatorial current $2I(\Psi_{\rm S})$ (notice that the equatorial current sheet contains double the current of each separatrix current sheet). Therefore, the only way to satisfy eq. (12) is to have $\left.B_{p}\right|_{\rm IN}\neq 0$. This can only be satisfied inside a T-point, not inside a Y-point. Nevertheless, we will continue calling the tip of the closed-line region the Y-point because of its overall bifurcating shape, and we will denote the rescaled radial distance of the tip of the closed-line region with $x_{\rm Y}$.
Furthermore, it is obvious that eq. (12) becomes singular when $x_{\rm Y}$ approaches very close to unity. In that limit, eq. (12) yields
$\left.B_{p}(x_{\rm Y})\right|_{\rm IN}=\frac{I(\Psi_{\rm S})}{x_{\rm Y}\sqrt{1-x_{\rm Y}^{2}}}=\frac{I(\Psi_{\rm S})}{\sqrt{2}\sqrt{1-x_{\rm Y}}}\rightarrow\infty\ \ \ \mbox{when}\ \ x_{\rm Y}\rightarrow 1\ .$ (13)
As is shown in this paper, contrary to what was obtained in previous works, it is not possible to reach this limit and have a significant open-line region outside. The closer the Y-point approaches the light cylinder, the more field lines must be squeezed inside the closed-line region to support the growth of $B_{p}$ at its tip required by eq. (13), the less field lines are left in the open-line region. As is shown with the new confined magnetospheric solution of figure 1, in the limit $x_{\rm Y}\rightarrow 1^{-}$ the open-line region disappears completely.

### 11 Apr 2024 v2

@@ 5 Discussion and Conclusions
We have improved the solutions of the pulsar equation when the tip of the closed-line region, the so-called Y-point, approaches the light cylinder. The new solutions were obtained under the assumption that the separatrix surface between open and closed field lines is a zero-thickness contact discontinuity. The new solutions are similar to the old ones of Contopoulos et al. (1999); Contopoulos (2005); Timokhin (2006) when the Y-point lies inside about $80\%$ of the light cylinder radius. As the Y-point approaches closer to the light cylinder, however, the solutions diverge from the standard ones found in the literature, and in the extreme limit that the closed-line region touches the light cylinder, the open-line region disappears completely. As is seen in Equation (12) in Appendix B and in Figure 4, the closer the Y-point to the light cylinder, the stronger the divergence of ${\rm B}_{p}$ right inside the Y-point at the tip of the closed-line region needs to be. We see no other way to implement this divergence of ${\rm B}_{p}(x_{\rm Y}\rightarrow 1^{-})$ except with more and more field lines to enter the closed-line region and compress the magnetic field at its tip. This effect is clearly seen in the sequence of solutions shown in Figure 3. In the limit where ${\rm B}_{p}\rightarrow\infty$ as $x_{\rm Y}=1^{-}$, all field lines enter the closed-line region confined inside the light cylinder, and the open line region disappears completely (Figure 1).
@@
One may legitimately ask why these solutions were never before obtained in the literature. There are several reasons. Firstly, in all previous solutions, the separatrix surface along which the magnetospheric electric current circuit closes on the star had non-zero thickness due to the finite resolution of the code. In particular, at its tip, the outer part of the separatrix forms a Y-point that touches the light cylinder, while the inner part forms a clear T-point some distance inside it. This is clearly explained in Figure 4 of Contopoulos et al. (2024). Unfortunately, the ideal force-free conditions of Equation (2) are not valid inside current sheets (as, e.g., in Figure 1b of Spitkovsky (2006)). Nevertheless, the fact that PIC simulations also form a separatrix that contains a nonzero amount of poloidal magnetic flux (see high-resolution detail in Figure 9 of Hu and Beloborodov (2022)) led us to consider the possibility that this effect is indeed physical. It is interesting that the separatrix current sheet contains a nonzero guide field and is thus different from the equatorial current sheet which is an ordinary Harris-type current sheet with $180^{\circ}$ field inversion across it. It also seems that the width of the former type of current sheet grows with time as is seen in Figure 1c of Spitkovsky (2006). This effect merits further investigation with hybrid PIC-FFE simulations (Soudais, Cerutti & Contopoulos 2024, submitted). Now that the separatrix has zero thickness, it forms a clear T-point at its tip, which indeed lies some distance inside the light cylinder. Secondly, time-dependent simulations always start from a vacuum dipole configuration that extends to all space and therefore, some fraction of it already crosses the light cylinder. After the central star is set into rotation, the magnetosphere is divided into open and closed field lines, i.e., there is no chance to obtain a configuration totally enclosed inside the light cylinder as the solution shown in Figure 1. Thirdly, after the time-dependent evolution relaxes to a steady-state solution, it is reasonable to expect that it will further relax to the lowest energy solution. Therefore, time-dependent simulations cannot yield the full sequence of solutions up to the latter singular solution as our Machine Learning methodology can. It is interesting that, now that the full sequence of solutions is obtained, it is shown that the solution where $x_{\rm Y}\sim 80\%\ R_{\rm LC}$ is a minimum energy solution, and this is why previous time-dependent high-resolution PIC simulations relax to that solution. We have thus improved upon the result of Contopoulos et al. (2024) which placed the minimum energy solution around $x_{\rm Y}=92\%\ R_{\rm LC}$.
@@
We argue that the confined singular solution of Figure 1 and all solutions between that one and the minimum energy solution are unstable; thus, if the magnetospheric configuration of Figure 1 forms during a supernova explosion or during a high accretion event from a surrounding disk or wind from a donor star, it will transition to the minimum energy solution by emitting a burst of energy equal to the energy difference between these two solutions. We propose that this burst may be related to an FRB. We would like to notice here that, as long as the surrounding material is there, the confined solution cannot transition to a standard solution in which part of the magnetosphere extends to infinity, and thus the accumulated electromagnetic energy of the confined solution cannot be released. What happens when the surrounding material dissolves and how fast electromagnetic energy is released remain to be seen in a time-dependent numerical simulation of this transition. In case the pulsar magnetosphere undergoes a later accretion event from an external disk or wind from a donor star that will re-confine it to the solution of Figure 1, the FRB will repeat. We must acknowledge of course that a few FRBs have already been observed from a magnetar, and that the cosmological sky rate of FRBs exceeds that of core-collapse supernovae; therefore, our model may just be a rare kind of FRB, if feasible at all. Moreover, we have not answered why would the emission of this state transition be in the MHz to GHz radio regime, and how it would escape the very dense accretion environment that formed the confined solution of Figure 1. Our model certainly merits further investigation with detailed MHD numerical simulations.
@@
What is also interesting to investigate is whether a confined solution exists for non-aligned (oblique) pulsars in which $\partial{\bf B}/\partial t\neq{\bf 0}$. We expect that oblique rotators would radiate electromagnetic radiation to infinity; thus, magnetic field lines would extend to infinity. Therefore, it seems impossible for an oblique magnetosphere to be totally confined inside the light cylinder. It may thus be that the confined solution of the type shown in Figure 1 is only possible in aligned pulsars.
@@
Conceptualization and supervision, I.C.; methodology, I.C. and K.N.G.; software and visualization, D.N. and I.D.; writing—original draft preparation, I.C.; writing—review and editing, all authors. All authors have read and agreed to the published version of the manuscript.
@@
This research work was partially funded by the Hellenic Foundation for Research and Innovation (HFRI).
@@ Acknowledgements.
Appendix A
In the first paper of this series (Dimitropoulos et al. (2024), Paper I), a new methodology was presented for solving the pulsar equation which essentially consists of the following steps:
@@
    It is very helpful to deal with the equatorial current sheet in the open line region by mathematically reversing the polarity of the magnetic field that emanates from the south pole of the star. By carrying this out, the mathematical discontinuity in the equatorial region disappears; thus, one does not need to worry about it computationally. At the end of the calculation, we will return to the original polarity of the magnetic field, and the equatorial current sheet will reappear.
@@
    The shape $r_{\rm S}(\theta)$ of the separatrix is readjusted so that continuity of ${\rm B}^{2}-{\rm E}^{2}$ is achieved at all points across it. One way to carry that out is to move the separatrix radially at each point with an amount that is proportional to the difference $({\rm B}^{2}-{\rm E}^{2})_{\rm IN}-({\rm B}^{2}-{\rm E}^{2})_{\rm OUT}$.
@@
    Repeat the procedure until pressure balance is achieved over the whole separatrix and its shape stabilizes.
@@
As it is acknowledged in Dimitropoulos et al. (2024), the numerical accuracy of our method is not on par yet with that of standard methods such as finite difference, finite volume or spectral methods. This is also the case in the solution of Stefanou et al. (2023). On the other hand, the numerical treatment of current sheets by standard methods it too is problematic (numerical current sheets have unphysical thicknesses and the numerical dissipation there is questionable). In order to improve our results in future iterations of this work, we will try to combine standard methods outside current sheets with our methodology of moving the separatrix. This is still a work in progress. Modulo the above disclaimers, we applied this method to obtain the sequence of solutions shown in Figure 3.
Appendix B
We discuss here the shape of the tip of the closed-line region, the so-called Y-point. In Paper I, it was found that
@@
    The Y-point is a clear T-point as predicted by Uzdensky (2003), not a Y-point as is seen in all numerical solutions of the axisymmetric pulsar magnetosphere to date, except for Contopoulos et al. (2024). This is probably due to the fact that in all previous numerical simulations, the separatrix between closed and open field lines had an unphysically large thickness (sometimes on the order of the light cylinder!); thus, it was not a true contact discontinuity as in Paper I.
@@
Both of the above results modify our understanding of how fast pulsars spin down. In the standard solution, the two regions are separated by a separatrix surface that contains a charged current sheet. One very interesting point in understanding the operation of the pulsar magnetosphere is how close to the light cylinder can the Y-point lie. According to Equation (3), the pressure balance condition across the separatrix current sheet may be written as
$\left.{\rm B}_{p}^{2}(1-x^{2})\right|_{\rm IN}=\left.{\rm B}_{p}^{2}(1-x^{2})\right|_{\rm OUT}+\frac{I(\Psi_{\rm S})^{2}}{x^{2}}$ (12)
Here, $I(\Psi_{\rm S})$ is the electric return current flowing along the separatrix, and $p$-indices denote poloidal ($r,\theta$) components of the fields. Let us first repeat the main argument why the tip of the closed line region on the equator is a T-point. The poloidal magnetic field $\left.{\rm B}_{p}\right|_{\rm OUT}$ in the right hand side of Equation (12) is obviously equal to zero right above and below the equatorial current sheet. The third term, however, is finite because of the presence of the nonzero equatorial current $2I(\Psi_{\rm S})$ (notice that the equatorial current sheet contains double the current of each separatrix current sheet). Therefore, the only way to satisfy Equation (12) is to have $\left.{\rm B}_{p}\right|_{\rm IN}\neq 0$. This can only be satisfied inside a T-point, not inside a Y-point. Nevertheless, we will continue calling the tip of the closed-line region the Y-point because of its overall bifurcating shape, and we will denote the rescaled radial distance of the tip of the closed-line region with $x_{\rm Y}$.
Furthermore, it is obvious that Equation (12) becomes singular when $x_{\rm Y}$ approaches very close to unity. In that limit, Equation (12) yields
$\left.{\rm B}_{p}(x_{\rm Y})\right|_{\rm IN}=\frac{I(\Psi_{\rm S})}{x_{\rm Y}\sqrt{1-x_{\rm Y}^{2}}}=\frac{I(\Psi_{\rm S})}{\sqrt{2}\sqrt{1-x_{\rm Y}}}\rightarrow\infty\ \ \ \mbox{when}\ \ x_{\rm Y}\rightarrow 1\ .$ (13)
As is shown in this paper, contrary to what was obtained in previous works, it is not possible to reach this limit and have a significant open-line region outside. The closer the Y-point approaches the light cylinder, the more field lines must be squeezed inside the closed-line region to support the growth of ${\rm B}_{p}$ at its tip required by Equation (13), the fewer field lines are left in the open-line region. As is shown with the new confined magnetospheric solution of Figure 1, in the limit $x_{\rm Y}\rightarrow 1^{-}$ the open-line region disappears completely.
[…]